# Orbital orders under magnetic fields in cubic PrIr2Zn20

### v3

@@ II.2 Landau Expansion
We now focus on the order parameter at the four equivalent L points. Note that ${\bm{k}}_{1}+{\bm{k}}_{2}+{\bm{k}}_{3}+{\bm{k}}_{4}\equiv{\bf 0}$ and that the fields at the L points are real since ${\bm{k}}_{n}\equiv-{\bm{k}}_{n}$.
The Landau’s free energy $F$ for the quadrupole moments is written as
$F$ $=\sum_{{\bm{k}}}[{\boldsymbol{\phi}}^{A}_{-{\bm{k}}},{\boldsymbol{\phi}}^{B}_{-{\bm{k}}}]\left[\frac{T}{2}+\hat{J}({\bm{k}})\right]\begin{bmatrix}{\boldsymbol{\phi}}^{A}_{{\bm{k}}}\\ {\boldsymbol{\phi}}^{B}_{{\bm{k}}}\end{bmatrix}$
@@
$+\frac{c}{4}\sum_{{\bm{k}},{\bm{k}}^{\prime},{\bm{k}}^{\prime\prime},{\bm{k}}^{\prime\prime\prime}}({\boldsymbol{\phi}}^{\alpha}_{{\bm{k}}}\cdot{\boldsymbol{\phi}}^{\alpha}_{{\bm{k}}^{\prime}})({\boldsymbol{\phi}}^{\alpha}_{{\bm{k}}^{\prime\prime}}\cdot{\boldsymbol{\phi}}^{\alpha}_{{\bm{k}}^{\prime\prime\prime}})\delta_{{\bm{k}}+{\bm{k}}^{\prime}+{\bm{k}}^{\prime\prime}+{\bm{k}}^{\prime\prime\prime},{\bm{G}}}\Big{]}.$ (5)
Here, $\hat{J}({\bm{k}})$ is the $4\times 4$ matrix in Eq. (2) with its eigenvalue $J({\bm{k}})$ and $T$ is the temperature and $\bm{G}$ is a reciprocal lattice vector. We have taken into account local cubic and fourth-order terms only, which arise from local anisotropy such as CEF potential.
@@ II.3 L point free energy
We now concentrate on the L point components ${\boldsymbol{\phi}}^{\alpha}_{n}=(u_{n}^{\alpha},v_{n}^{\alpha})\equiv{\boldsymbol{\phi}}^{\alpha}_{{\bm{k}}_{n}}(n=1,2,3,4)$ with the short handed notation, we find the part of the free energy consisting of only ${\boldsymbol{\phi}}^{\alpha}_{{\bm{k}}_{n}}$’s is
$F_{\rm L}$ $=\frac{a_{\rm L}}{2}\sum_{\alpha,n}|{\boldsymbol{\phi}}_{n}^{\alpha}|^{2}+J_{\rm L}\sum_{n}(-1)^{p_{n}}{\boldsymbol{\phi}}^{A}_{n}\cdot{\boldsymbol{\phi}}^{B}_{n}+\frac{c}{4}\sum_{\alpha,n}|{\boldsymbol{\phi}}_{n}^{\alpha}|^{4}$
$+\frac{c^{\prime}}{4}\sum_{\alpha}\sum_{m<n}|{\boldsymbol{\phi}}^{\alpha}_{m}|^{2}|{\boldsymbol{\phi}}^{\alpha}_{n}|^{2}+\frac{c^{\prime\prime}}{4}\sum_{\alpha}\sum_{m<n}({\boldsymbol{\phi}}_{m}^{\alpha}\cdot{\boldsymbol{\phi}}_{n}^{\alpha})^{2}$
$+\frac{c^{\prime\prime\prime}}{4}\sum_{\alpha}\sum_{(m,n,m^{\prime},n^{\prime})}({\boldsymbol{\phi}}^{\alpha}_{m}\cdot{\boldsymbol{\phi}}^{\alpha}_{n})({\boldsymbol{\phi}}^{\alpha}_{m^{\prime}}\cdot{\boldsymbol{\phi}}^{\alpha}_{n^{\prime}}).$ (6)
Note that $b$ terms are absent owing to the wavnumber conservation and $(m,n,m^{\prime},n^{\prime})=(1,2,3,4)$, $(1,3,2,4)$, and $(1,4,2,3)$ in the last term in the right-hand side. The factor $p_{n}(n=1,2,3,4)$ represents the sign of exchange interaction between the $A$ and $B$ sublattices for four kinds of the L points. Since our unit cell consists of the bond parallel to ${\bm{k}}_{1}$, $p_{n}=0$ is distinct from the others $p_{2,3,4}=1$. The coupling constants are denoted as
@@
Note that ${\bm{k}}_{n}\equiv-{\bm{k}}_{n}$ and thus the ${\boldsymbol{\phi}}_{n}\cdot{\boldsymbol{\phi}}_{n}$ pair in the summation in $\sum_{{\bm{k}}}$ appears once. Since each of the fourth order terms are allowed by the cubic symmetry, we denote them as independent parameters.
For discussing a single-$\bm{q}$ order at the L points, $c^{\prime}$, $c^{\prime\prime}$, and $c^{\prime\prime\prime}$ terms are irrelevant and the fourth-order free energy is simply given by
$F_{L}^{1\bm{q}(4)}=\frac{c}{4}\sum_{\alpha}|{\boldsymbol{\phi}}^{\alpha}|^{4}.$ (9)
When multiple-$\bm{q}$ states are considered, the coupling constants $c^{\prime}$, $c^{\prime\prime}$, and $c^{\prime\prime\prime}$ matter as will be discussed in Sec. V.
@@ II.4 $\Gamma$ point free energy
Under finite magnetic fields, the uniform quadrupole moment ${\boldsymbol{\phi}}_{0}^{\alpha}\equiv{\boldsymbol{\phi}}_{{\bm{k}}=\bf 0}^{\alpha}$ is induced via the effective field $\tilde{\bm{h}}$ in the second-order perturbation of the magnetic field $\bm{H}$. The free energy up to the fourth-order terms for the $\Gamma$ point modes are
$F_{\Gamma}$ $=\sum_{\alpha}\Big{(}-\tilde{\bm{h}}\cdot{\boldsymbol{\phi}}_{0}^{\alpha}+\frac{a_{\Gamma 0}}{2}|{\boldsymbol{\phi}}_{0}^{\alpha}|^{2}\Big{)}+J_{\Gamma 0}{\boldsymbol{\phi}}_{0}^{A}\cdot{\boldsymbol{\phi}}_{0}^{B}$
$+\sum_{\alpha}\Big{\{}-bu_{0}^{\alpha}\Big{[}\big{(}u_{0}^{\alpha}\big{)}^{2}-3\big{(}v_{0}^{\alpha}\big{)}^{2}\Big{]}+\frac{c}{4}|{\boldsymbol{\phi}}_{0}^{\alpha}|^{4}\Big{\}},$ (10)
@@
and $g=(g_{J}\mu_{\rm B})^{2}[7/(3E_{4})-1/E_{5}]$ with the excited $\Gamma_{4,5}$ CEF energy $E_{4,5}$.Hattori and Tsunetsugu (2014) $\mu_{\rm B}$ is the Bohr magneton and $g_{J}$ is the Lande’s $g$ factor for the multiplet with the angular momentum $J$; $g_{J}=4/5$ for the PrIr2Zn20 with $J=4$.
@@ II.5 Mode coupling terms
We now discuss mode coupling terms consisting of the fields at the L points and those at the $\Gamma$ points. The lowest-order contributions come from the cubic terms as
$\delta F_{3}=$ $-3\lambda_{b}\sum_{\alpha,n}\Big{\{}u_{0}^{\alpha}\Big{[}\big{(}u_{n}^{\alpha}\big{)}^{2}-\big{(}v_{n}^{\alpha}\big{)}^{2}\Big{]}-v_{0}^{\alpha}(2u_{n}^{\alpha}v_{n}^{\alpha})\Big{\}}$
@@
${\boldsymbol{\phi}}^{\alpha}_{\bf 0}$ $=\phi^{\alpha}_{0}\begin{bmatrix}\cos\theta^{\alpha}_{0}\\ \sin\theta^{\alpha}_{0}\end{bmatrix},\quad{\boldsymbol{\phi}}^{\alpha}_{n}=\phi^{\alpha}_{n}\begin{bmatrix}\cos\theta^{\alpha}_{n}\\ \sin\theta^{\alpha}_{n}\end{bmatrix},$ (14)
@@
$\delta F_{4}=\sum_{\alpha,n}\left[\frac{\lambda^{(1)}_{c}}{2}|{\boldsymbol{\phi}}_{n}^{\alpha}|^{2}|{\boldsymbol{\phi}}_{0}^{\alpha}|^{2}+\lambda_{c}^{(2)}({\boldsymbol{\phi}}_{n}^{\alpha}\cdot{\boldsymbol{\phi}}_{0}^{\alpha})^{2}\right].$ (15)
@@
In principle, ${\boldsymbol{\phi}}_{0}^{\alpha}$ can couple with ${\boldsymbol{\phi}}_{{\bm{k}}}^{\alpha}$ in the form of $\sim{\boldsymbol{\phi}}_{0}^{\alpha}{\boldsymbol{\phi}}_{\bm{k}}^{\alpha}{\boldsymbol{\phi}}_{-{\bm{k}}}^{\alpha}$ and ${\boldsymbol{\phi}}_{0}^{\alpha}{\boldsymbol{\phi}}_{0}^{\alpha}{\boldsymbol{\phi}}_{\bm{k}}^{\alpha}{\boldsymbol{\phi}}_{-{\bm{k}}}^{\alpha}$. In this paper, we retain the order parameter ${\boldsymbol{\phi}}_{n}^{\alpha}$ and the induced moment ${\boldsymbol{\phi}}_{0}^{\alpha}$ under magnetic fields since the other fields are not important for the phase transition with the L point fields. Some renormalization effects arising from fields at the X points will be discussed in Sec. IV and Appendix B.
@@ III Results: single-q orders
In this section, we carry out minimization of the free energy $F_{L}+F_{\Gamma}+\delta F_{3}+\delta F_{4}$ [Eqs. (6), (10), (13), and (15)], assuming the single-$\bm{q}$ orders. The purpose here is to understand how mode coupling effects affect the order parameter configuration. The results in this section will become bases for understanding the full phase diagram shown in Sec. IV, where we will take into account multiple-$\bm{q}$ configurations. Since we are interested in the phase transition owing to the order parameters at the L point, we assume the coefficients of the quadratic terms in the free energy at the other wavenumber are always positive. We first discuss the candidates of the stable configurations by analyzing the third- and fourth-order terms in $F_{\rm L}$. Then, we will show the numerical results of full temperature-field phase diagram.
@@ III.1 Qualitative analysis
First, we notice that the inter-sublattice correlations are governed by $J_{\rm L}=-2J_{1}+6J_{3}$ for ${\boldsymbol{\phi}}_{n}^{\alpha}$ in Eq. (6). To reproduce the quadrupole configuration propposed by the neutron scattering experimentsIwasa et al. (2017), one needs $J_{\rm L}<0$. Thus, it is legitimate to set
${\boldsymbol{\phi}}_{\rm L}\equiv{\boldsymbol{\phi}}_{1}^{A}={\boldsymbol{\phi}}_{1}^{B}=\begin{bmatrix}u_{\rm L}\\ v_{\rm L}\end{bmatrix}=\phi_{\rm L}\begin{bmatrix}\cos\theta_{\rm L}\\ \sin\theta_{\rm L}\end{bmatrix}.$ (16)
Here, we have chosen one of the L points at ${\bm{k}}_{1}$ as a representative one since the other three fields at ${\bm{k}}_{n}(n=2,3,4)$ with ${\boldsymbol{\phi}}_{n}^{A}=-{\boldsymbol{\phi}}_{n}^{B}$ are equivalent and this is also the case even when $\bm{H}$ is finite. See Table 1. For the modes at the $\Gamma$ point, which are gapped and just induced by magnetic fields or by the square of the order parameter at the L point, we set
${\boldsymbol{\phi}}_{0}\equiv{\boldsymbol{\phi}}_{0}^{A}={\boldsymbol{\phi}}_{0}^{B}=\begin{bmatrix}u_{0}\\ v_{0}\end{bmatrix}=\phi_{0}\begin{bmatrix}\cos\theta_{0}\\ \sin\theta_{0}\end{bmatrix}.$ (17)
Second, we note that the single-$\bm{q}$ free energy $F_{\rm L}$ is isotropic. This means that the transition temperatures for $u_{\rm L}$ and $v_{\rm L}$ are the same at ${\bm{H}}={\bf 0}$ and $T_{c}=6J_{4}+2|J_{1}-3J_{3}|$ This isotropic nature, however, is destroyed by the mode coupling to the fields at the $\Gamma$ points through $\delta F_{3}$ and $\delta F_{4}$. When $T\lesssim T_{c}$, the induced ${\boldsymbol{\phi}}_{0}$ is sufficiently small, and thus, the third- and fourth-order terms in $F_{\Gamma}$ can be irrelevant. In this situation the relative directions between ${\boldsymbol{\phi}}_{n}$ and ${\boldsymbol{\phi}}_{0}$ is determined by $\delta F_{3}$ and $\delta F_{4}$ as
@@
with $j=1,2,3$. These six domains are degenerate at ${\bm{H}}={\bf 0}\ (\tilde{\bm{h}}={\bf 0})$. See Fig. 2(a). We should comment that the three directions of ${\boldsymbol{\phi}}_{0}$ do not coincide with those favored by the $b(>0)$ term in $F_{\Gamma}:-b\phi_{0}^{3}\cos(3\theta_{0})$. Since ${\boldsymbol{\phi}}_{0}$ is not the primary order parameter, the third-order term is expected to play a minor role and the configurations (19) are robust at least $\tilde{\bm{h}}={\bf 0}$.
Figure 2: Schematic single-$\bm{q}$ order parameter configurations for (a) ${\bm{H}=0}$, (b) small ${\bm{H}}$, (c) small ${\bm{H}}$ and low $T$, and (d) large ${\bm{H}}$. Blue (red) arrows indicate ${\boldsymbol{\phi}}_{0}$(${\boldsymbol{\phi}}_{\rm L}$). Filled red circles represent the directions $\theta_{\rm L}$ given by Eq. (19) and open blue circles do $\theta_{\rm 0}$. A configuration for large ${\bm{H}}$ in the phase II is shown in (e) and that for ${\bm{H}}\parallel[110]$ in (f).
Now let us consider the cases for finite magnetic fields. For high symmetric directions $\bm{H}=\tfrac{H}{\sqrt{2}}(\sin\theta,\sin\theta,\sqrt{2}\cos\theta)$ with $20\leq\theta\leq\pi/$ 2, the effective field reads as $\tilde{h}_{u}=gH^{2}(3\cos^{2}\theta-1)$ and $\tilde{h}_{v}=0$. Note that $\tilde{\bm{h}}={\bf 0}$ at $\bm{H}\parallel[111]$ and $\tilde{h}_{u}>0$ for $\theta<\tan^{-1}\sqrt{2}\equiv\theta_{111}$, while $\tilde{h}_{u}<0$ for $\theta>\theta_{111}$. The effective field induces $u_{0}$, which leads to two “transition temperatures” $\tilde{T}_{c}^{u}$ for the $u_{\rm L}$ order and $\tilde{T}_{c}^{v}$ for $v_{\rm L}$. They are given by
@@
As is clear from the expression of $\Delta\tilde{T}_{c}$, it is expected that there is no additional phase transition within the single-$\bm{q}$ ${\boldsymbol{\phi}}_{n}$ sector for $u_{0}<0$, i.e., $\theta_{0}>\theta_{111}$. In the following, we will concentrate on the cases for $u_{0}>0$ ($\theta_{0}<\theta_{111}$). For $u_{0}>0$, the linear coupling between ${\boldsymbol{\phi}}_{0}$ and $\tilde{h}_{u}$ can alter the stable configuration from the above six domains (19) since their energy gain is given by the second order in ${\boldsymbol{\phi}}_{0}$. Since $\Delta\tilde{T}_{c}>0$, the phase transition from the disordered phase occurs in the $u_{\rm L}$ component: $\theta_{n}=0$ or $\pi$. Here, the two choices 00 and $\pi$ represent a simple translation and we will consider $\theta_{\rm L}=0$ below. For small $\tilde{h}_{u}$, $\delta F_{3}$ is more important than $\delta F_{4}$. This means $\theta_{0}=0$ satisfying $02\theta_{\rm L}+\theta_{0}=$ 0 [Fig. 2(b)]. Such a collinear state possesses the free energy $F_{\rm col}$ and the fluctuations around this is given as
$\delta F_{\rm col}$ $\simeq\lambda_{c}\phi_{0}\phi_{\rm L}^{2}\begin{bmatrix}\theta_{0}&\theta_{\rm L}\end{bmatrix}\hat{A}\begin{bmatrix}\theta_{0}\\ \theta_{\rm L}\end{bmatrix},$ (21)

### v4

@@ II.2 Landau Expansion
We focus on the order parameter at the four equivalent L points. Note that ${\bm{k}}_{1}+{\bm{k}}_{2}+{\bm{k}}_{3}+{\bm{k}}_{4}\equiv{\bf 0}$ and that the fields at the L points are real since ${\bm{k}}_{n}\equiv-{\bm{k}}_{n}$.
The Landau’s free energy $\mathcal{F}$} for the quadrupole moments is written as
${\mathcal{F}}$ $=\sum_{{\bm{k}}}[{\bm{\phi}}^{A}_{-{\bm{k}}},{\bm{\phi}}^{B}_{-{\bm{k}}}]\left[\frac{T}{2}+\hat{J}({\bm{k}})\right]\begin{bmatrix}{\bm{\phi}}^{A}_{{\bm{k}}}\\ {\bm{\phi}}^{B}_{{\bm{k}}}\end{bmatrix}$
@@
$+\frac{c}{4}\sum_{{\bm{k}},{\bm{k}}^{\prime},{\bm{k}}^{\prime\prime},{\bm{k}}^{\prime\prime\prime}}({\bm{\phi}}^{\alpha}_{{\bm{k}}}\cdot{\bm{\phi}}^{\alpha}_{{\bm{k}}^{\prime}})({\bm{\phi}}^{\alpha}_{{\bm{k}}^{\prime\prime}}\cdot{\bm{\phi}}^{\alpha}_{{\bm{k}}^{\prime\prime\prime}})\delta_{{\bm{k}}+{\bm{k}}^{\prime}+{\bm{k}}^{\prime\prime}+{\bm{k}}^{\prime\prime\prime},{\bm{G}}}\Big].$ (5)
Here, $\hat{J}({\bm{k}})$ is the $4\times 4$ matrix in Eq. (2) with its eigenvalue $J({\bm{k}})$, $T$ is the temperature, and $\bm{G}$ is a reciprocal lattice vector. We have taken into account local cubic and fourth-order terms only, which arise from local anisotropy such as CEF potential. We note that the cubic term $b$ arises from the effects of the excited states and relatively smaller than $c$ as estimated by using the actual CEF scheme [21].
@@ II.3 L point free energy
We concentrate on the L point components ${\bm{\phi}}^{\alpha}_{n}=(u_{n}^{\alpha},v_{n}^{\alpha})\equiv{\bm{\phi}}^{\alpha}_{{\bm{k}}_{n}}(n=1,2,3,4)$ with the short-handed notation, we find the part of the free energy consisting of only ${\bm{\phi}}^{\alpha}_{{\bm{k}}_{n}}$’s is
${\mathcal{F}}_{\rm L}$ $=\frac{a_{\rm L}}{2}\sum_{\alpha,n}|{\bm{\phi}}_{n}^{\alpha}|^{2}+J_{\rm L}\sum_{n}(-1)^{p_{n}}{\bm{\phi}}^{A}_{n}\cdot{\bm{\phi}}^{B}_{n}+\frac{c}{4}\sum_{\alpha,n}|{\bm{\phi}}_{n}^{\alpha}|^{4}$
$+\frac{c^{\prime}}{4}\sum_{\alpha}\sum_{m<n}|{\bm{\phi}}^{\alpha}_{m}|^{2}|{\bm{\phi}}^{\alpha}_{n}|^{2}+\frac{c^{\prime\prime}}{4}\sum_{\alpha}\sum_{m<n}({\bm{\phi}}_{m}^{\alpha}\cdot{\bm{\phi}}_{n}^{\alpha})^{2}$
$+\frac{c^{\prime\prime\prime}}{4}\sum_{\alpha}\sum_{(m,n,m^{\prime},n^{\prime})}({\bm{\phi}}^{\alpha}_{m}\cdot{\bm{\phi}}^{\alpha}_{n})({\bm{\phi}}^{\alpha}_{m^{\prime}}\cdot{\bm{\phi}}^{\alpha}_{n^{\prime}}).$ (6)
Note that $b$ terms are absent owing to the wavenumber conservation and $(m,n,m^{\prime},n^{\prime})=(1,2,3,4)$, $(1,3,2,4)$, and $(1,4,2,3)$ in the last term in the right-hand side. The factor $p_{n}(n=1,2,3,4)$ represents the sign of exchange interaction between the $A$ and $B$ sublattices for four types of L points. Since our unit cell consists of the bond parallel to ${\bm{k}}_{1}$, $p_{n}=0$ is distinct from the others $p_{2,3,4}=1$. The coupling constants are denoted as
@@
Note that ${\bm{k}}_{n}\equiv-{\bm{k}}_{n}$ and thus the ${\bm{\phi}}_{n}\cdot{\bm{\phi}}_{n}$ pair in the summation in $\sum_{{\bm{k}}}$ appears once. Since the three fourth-order terms are independently allowed by the cubic symmetry, we denote them as independent parameters.
For discussing a single-$\bm{q}$ order at the L points, $^{\prime}, c′′c^{\prime\prime}$, and $c^{\prime\prime\prime}$ terms are irrelevant and the fourth-order free energy is simply given by
${\mathcal{F}}_{L}^{1\bm{q}(4)}=\frac{c}{4}\sum_{\alpha}|{\bm{\phi}}^{\alpha}|^{4}.$ (9)
When multiple-$\bm{q}$ states are considered, the coupling constants $^{\prime}, c′′c^{\prime\prime}$, and $c^{\prime\prime\prime}$ matter as discussed in Sec. V.
@@ II.4 $\Gamma$ point free energy
Under finite magnetic fields, the uniform quadrupole moment ${\bm{\phi}}_{0}^{\alpha}\equiv{\bm{\phi}}_{{\bm{k}}=\bf 0}^{\alpha}$ is induced via the effective field $\tilde{\bm{h}}$ in the second-order perturbation of the magnetic field $\bm{H}$. The free energy up to the fourth-order terms for the $\Gamma$-point modes are
${\mathcal{F}}_{\Gamma}$ $=\sum_{\alpha}\Big(-\tilde{\bm{h}}\cdot{\bm{\phi}}_{0}^{\alpha}+\frac{a_{\Gamma 0}}{2}|{\bm{\phi}}_{0}^{\alpha}|^{2}\Big)+J_{\Gamma 0}{\bm{\phi}}_{0}^{A}\cdot{\bm{\phi}}_{0}^{B}$
$+\sum_{\alpha}\Big\{-bu_{0}^{\alpha}\Big[\big(u_{0}^{\alpha}\big)^{2}-3\big(v_{0}^{\alpha}\big)^{2}\Big]+\frac{c}{4}|{\bm{\phi}}_{0}^{\alpha}|^{4}\Big\},$ (10)
@@
and $g=(g_{J}\mu_{\rm B})^{2}[7/(3E_{4})-1/E_{5}]$ with the excited $\Gamma_{4,5}$ CEF energy $E_{4,5}$ [21]. $\mu_{\rm B}$ is the Bohr magneton and $g_{J}$ is the Lande’s $g$ factor for the multiplet with the angular momentum $J$; $g_{J}=4/5$ for the PrIr2Zn20 with $J=4$.
@@ II.5 Mode coupling terms
We now discuss mode coupling terms consisting of the fields at the L points and those at the $\Gamma$ point. The lowest-order contributions come from the cubic terms as
$\delta{\mathcal{F}}_{3}=$ $-3\lambda_{b}\sum_{\alpha,n}\Big\{u_{0}^{\alpha}\Big[\big(u_{n}^{\alpha}\big)^{2}-\big(v_{n}^{\alpha}\big)^{2}\Big]-v_{0}^{\alpha}(2u_{n}^{\alpha}v_{n}^{\alpha})\Big\}$
@@
${\bm{\phi}}^{\alpha}_{\bf 0}$ $=\phi^{\alpha}_{0}\begin{bmatrix}\cos\theta^{\alpha}_{0}\\ \sin\theta^{\alpha}_{0}\end{bmatrix},\quad{\bm{\phi}}^{\alpha}_{n}=\phi^{\alpha}_{n}\begin{bmatrix}\cos\theta^{\alpha}_{n}\\ \sin\theta^{\alpha}_{n}\end{bmatrix}$, (14)
@@
$\delta{\mathcal{F}}_{4}=\sum_{\alpha,n}\left[\frac{\lambda^{(1)}_{c}}{2}|{\bm{\phi}}_{n}^{\alpha}|^{2}|{\bm{\phi}}_{0}^{\alpha}|^{2}+\lambda_{c}^{(2)}({\bm{\phi}}_{n}^{\alpha}\cdot{\bm{\phi}}_{0}^{\alpha})^{2}\right].$ (15)
@@
In principle, ${\bm{\phi}}_{0}^{\alpha}$ can couple with ${\bm{\phi}}_{{\bm{k}}}^{\alpha}$ in the form of $\sim{\bm{\phi}}_{0}^{\alpha}{\bm{\phi}}_{\bm{k}}^{\alpha}{\bm{\phi}}_{-{\bm{k}}}^{\alpha}$ and ${\bm{\phi}}_{0}^{\alpha}{\bm{\phi}}_{0}^{\alpha}{\bm{\phi}}_{\bm{k}}^{\alpha}{\bm{\phi}}_{-{\bm{k}}}^{\alpha}$. In this paper, we retain the order parameter ${\bm{\phi}}_{n}^{\alpha}$ and the induced moment ${\bm{\phi}}_{0}^{\alpha}$ under magnetic fields since the other fields are not important for the phase transition with the L point fields. Some renormalization effects arising from fields at the X points are discussed in Sec. IV and Appendix B.
@@ III Results: single-q orders
In this section, we carry out minimization of the free energy ${\mathcal{F}}_{L}+{\mathcal{F}}_{\Gamma}+\delta{\mathcal{F}}_{3}+\delta{\mathcal{F}}_{4}$ [Eqs. (6), (10), (13), and (15)], assuming the single-$\bm{q}$ orders. The purpose here is to understand how mode coupling effects affect the order parameter configuration. The results in this section will serve as the basis for understanding the full phase diagram shown in Sec. IV, where we will take into account multiple-$\bm{q}$ configurations. Since we are interested in the phase transition owing to the order parameters at the L point, we assume the coefficients of the quadratic terms in the free energy at ${\bm{k}}\neq{\bm{k}}_{n}$ are always positive. We first discuss the candidates of the stable configurations by analyzing the third- and fourth-order terms in ${\mathcal{F}}_{\rm L}$. Then, we will show the numerical results of full temperature-field phase diagram.
@@ III.1 Qualitative analysis
First, we notice that the inter-sublattice correlations are governed by $J_{\rm L}=-2J_{1}+6J_{3}$ for ${\bm{\phi}}_{n}^{\alpha}$ in Eq. (6). To reproduce the quadrupole configuration propposed by the neutron scattering experiments [34], $J_{\rm L}<0$ is required. Thus, it is legitimate to set
${\bm{\phi}}_{\rm L}\equiv{\bm{\phi}}_{1}^{A}={\bm{\phi}}_{1}^{B}=\begin{bmatrix}u_{\rm L}\\ v_{\rm L}\end{bmatrix}=\phi_{\rm L}\begin{bmatrix}\cos\theta_{\rm L}\\ \sin\theta_{\rm L}\end{bmatrix}$. (16)
Here, we have chosen one of the L points at ${\bm{k}}_{1}$ as a representative one since the other three fields at ${\bm{k}}_{n}(n=2,3,4)$ with ${\bm{\phi}}_{n}^{A}=-{\bm{\phi}}_{n}^{B}$ are equivalent and this is also the case even when $\bm{H}$ is finite. See Table 1. For the modes at the $\Gamma$ point, which are gapped and just induced by magnetic fields or by the square of the order parameter at the L point, we set
${\bm{\phi}}_{0}\equiv{\bm{\phi}}_{0}^{A}={\bm{\phi}}_{0}^{B}=\begin{bmatrix}u_{0}\\ v_{0}\end{bmatrix}=\phi_{0}\begin{bmatrix}\cos\theta_{0}\\ \sin\theta_{0}\end{bmatrix}$. (17)
Second, we note that the single-$\bm{q}$ free energy ${\mathcal{F}}_{\rm L}$ is isotropic. This means that the transition temperatures for $u_{\rm L}$ and $v_{\rm L}$ are the same at ${\bm{H}}={\bf 0}$ and $T_{c}=6J_{4}+2|J_{1}-3J_{3}|$ . This isotropic nature, however, is destroyed by the mode coupling to the fields at the $\Gamma$ point through $\delta{\mathcal{F}}_{3}$ and $\delta{\mathcal{F}}_{4}$. When $T\lesssim T_{c}$, the induced ${\bm{\phi}}_{0}$ is sufficiently small, and thus, the third- and fourth-order terms in ${\mathcal{F}}_{\Gamma}$ can be irrelevant. In this situation the relative directions between ${\bm{\phi}}_{n}$ and ${\bm{\phi}}_{0}$ is determined by $\delta{\mathcal{F}}_{3}$ and $\delta{\mathcal{F}}_{4}$ as
@@
with $j=1,2,3$. These six domains are degenerate at ${\bm{H}}={\bf 0}\ (\tilde{\bm{h}}={\bf 0})$. See Fig. 3(a). We should comment that the three directions of ${\bm{\phi}}_{0}$ do not coincide with those favored by the $b(>0)$ term in ${\mathcal{F}}_{\Gamma}:-b\phi_{0}^{3}\cos(3\theta_{0})$. Since ${\bm{\phi}}_{0}$ is not the primary order parameter, the third-order term is expected to play a minor role and the configurations (19) are robust at least $\tilde{\bm{h}}={\bf 0}$.
Figure 3: Schematic single-$\bm{q}$ order parameter configurations for (a) $\bm{H}=0$}, (b) small $\bm{H}$}, (c) small $\bm{H}$} and low $T$, and (d) large $\bm{H}$}. Blue (red) arrows indicate ${\bm{\phi}}_{0}$(${\bm{\phi}}_{\rm L}$). Filled red circles represent the directions $\theta_{\rm L}$ given by Eq. (19) and open blue circles do $\theta_{\rm 0}$. A configuration for large $\bm{H}$} in the phase II is shown in (e) and that for ${\bm{H}}\parallel[110]$ in (f).
Figure 4: (a)–(c) TTHH phase diagram for $\bm{H}\parallel[001]$, $J_{1}=1.0$, $J_{2}=J_{3}=0.0$, $J_{4}=1.0$, and $c=0.5$. (a) $b=0.1$, (b) $b=0.05$, and (c) $b=0.01$. Color map represents the L point order parameter angle $\theta_{\rm L}$. (d) TTHH phase diagram for $\bm{H}\parallel[110]$ with the same parameter set as in (b). The schematic L point order parameter configuration for (e) the phase I and for (f) the phases III and IV. For clarity, orbital shapes are drawn only for four sites. Red (white) spheres represent Pr sites with $\theta_{\rm L}=0\ (\pi)$ for (e) and $\theta_{\rm L}=\pi/2\ (-\pi/2)$ for (f), respectively. On shaded (111) planes, the order parameter $\bm{\phi}^{A}(\bm{r})$ is uniform and $\bm{\phi}^{B}(\bm{r})=\bm{\phi}^{A}(\bm{r})$. For the phase II, the L point order parameter is in between those in the phase I and III with $0<|\theta_{\rm L}|<\pi/2$.
Now let us consider the cases for finite magnetic fields. For high symmetric directions $\bm{H}=\tfrac{H}{\sqrt{2}}(\sin\Theta,\sin\Theta,\sqrt{2}\cos\Theta)$ with $0\leq\Theta\leq\pi/2$, the effective field reads as $\tilde{h}_{u}=gH^{2}(3\cos^{2}\Theta-1)$ and $\tilde{h}_{v}=0$. Note that $\tilde{\bm{h}}={\bf 0}$ at $\bm{H}\parallel[111]$ and $\tilde{h}_{u}>0$ for $\Theta<\tan^{-1}\sqrt{2}\equiv\Theta_{111}$, while $\tilde{h}_{u}<0$ for $\Theta>\Theta_{111}$. The effective field induces $u_{0}$, which leads to two “transition temperatures” $\tilde{T}_{c}^{u}$ for the $u_{\rm L}$ order and $\tilde{T}_{c}^{v}$ for $v_{\rm L}$. They are given by
@@
As is clear from the expression of $\Delta\tilde{T}_{c}$, it is expected that there is no additional phase transition within the single-$\bm{q}$ ${\bm{\phi}}_{n}$ sector for $u_{0}<0$, i.e., $\Theta>\Theta_{111}$. In the following, we will concentrate on the cases for $u_{0}>0$ ($\Theta<\Theta_{111}$). For $u_{0}>0$, the linear coupling between ${\bm{\phi}}_{0}$ and $\tilde{h}_{u}$ can alter the stable configuration from the above six domains (19) since their energy gain is given by the second order in ${\bm{\phi}}_{0}$. Since $\Delta\tilde{T}_{c}>0$, the phase transition from the disordered phase occurs in the $u_{\rm L}$ component: $\theta_{\rm L}=0$ or $\pi$. Here, the two choices 0 and $\pi$ represent a simple translation and we will consider $\theta_{\rm L}=0$ below. For small $\tilde{h}_{u}$, $\delta{\mathcal{F}}_{3}$ is more important than $\delta{\mathcal{F}}_{4}$. This means $\theta_{0}=0$ satisfying $2\theta_{\rm L}+\theta_{0}=0$ [Fig. 3(b)]. Such a collinear state possesses the free energy ${\mathcal{F}}_{\rm col}$ and the fluctuations around this is given as
$\delta{\mathcal{F}}_{\rm col}$ $\simeq\lambda_{c}\phi_{0}\phi_{\rm L}^{2}\begin{bmatrix}\theta_{0}&\theta_{\rm L}\end{bmatrix}\hat{A}\begin{bmatrix}\theta_{0}\\ \theta_{\rm L}\end{bmatrix}$, (21)